Exercise 11-12 Volume Trade-Off Decisions [LO11-5] Benoit Company produces three products-A, B, and C. Data concerning the three products follow (per unit Product Selling price Variable expenses: 586.6 $66.80 $76.6 Direct materials Other variable expenses Total variable expenses Contribution margin Contribution margin ratio 25.88 18.00 9.00 25.80 31.5 44.20 51.60 49.5 53.20 534.40 $16.50 $22.80 40% 30% The company estimates that it can sell 950 units of each product per month. The same raw material is used in each product. The material costs $3 per pound with a maximum of 5.000 pounds available each month. Required 1. Calculate the contribution margin per pound of the constraining resource for each product 2. Which orders would you advise the company to accept first, those for A, B, or C? Which or 3. What is the maximum cont tribution margin that the company can earn per month if it makes optimal use of its 5,000 pounds of Complete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Calculate the contribution margin per pound of the constraining resource for each product.(Round your answers to 2 decimal places.) Product A Product B Product C Contribution margin per pound of the constraining resource Required 2>
Answer #1

Answers

  • All working forms part of the answer
  • Requirement 1

Working

A

B

C

A

Directy Materials per unit

$                   25.80

$                  18.00

$                    9.00

B

Material cost per pound

$                     3.00

$                    3.00

$                    3.00

C = A / B

Direct material pound per unit

$                     8.60

6

3

D

Contribution margin per unit

$                   34.40

$                  16.50

$                  22.80

E = D/C

Contribution margin per pound [ ANSWER]

$                     4.00

$                    2.75

$                    7.60

  • Requirement 2

Order will be based on Highest Contribution margin per pound to Lowest

Product A

SECOND

Product B

THIRD

Product C

FIRST

  • Requirement 3

Answer = 8600 + 21660 = $ 30,260 Maximum Contribution

----Working

A

B

C

Maximum demands

                          950

                         950

                         950

Direct material pound per unit

                        8.60

                        6.00

                        3.00

Total Material requirement for fulfulling demand

                      8,170

                     5,700

                     2,850

Total pounds Available

A

B

C

pounds Allocated

Total Material requirement for fulfulling demand

                     8,170

                     5,700

                2,850

Pounds available

                      5,000

#1 Allocation to C

                2,850

                    2,850

Pounds available

                      2,150

#2 Allocation to A

                     2,150

Pounds available

                             -  

Total pounds of material allocated

                     2,150

                            -  

                2,850

Contribution margin per pound

$                    4.00

$                    2.75

$               7.60

Total Contribution margin

$            8,600.00

$                         -  

$    21,660.00